OpenStudy (anonymous):

the hypotenuse of a right triangle is 3 less than twice the shorter leg. The length of the other leg is 3 more than the shorter leg. Find the length of the shorter leg. (quadratic equations)

OpenStudy (anonymous):

\[h=2a-3\] \[b=a+3\] \[h^2=a^2+b^2\]

OpenStudy (anonymous):

like they say quadratic ,put what h and b is , in the quadratic eq

OpenStudy (anonymous):

they meant that we have to use quadratic eq to solvefor the smallest side

OpenStudy (anonymous):

ooohhhh, wait

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get it now!!!

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

you substitute h and b into the third eq, right?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so, (2a-3)^2=a2 + (a+3)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes then 2a^2-18a=0 a=0 ir a=9